# Unified World Models: Coupling Video and Action Diffusion for Pretraining on Large Robotic Datasets

### v2

@@ Abstract
Imitation learning has emerged as a promising approach towards building generalist robots. However, scaling imitation learning for large robot foundation models remains challenging due to its reliance on high-quality expert demonstrations. Meanwhile, large amounts of video data depicting a wide range of environments and diverse behaviors are readily available. This data provides a rich source of information about real-world dynamics and agent-environment interactions. Leveraging this data directly for imitation learning, however, has proven difficult due to the lack of action annotation required for most contemporary methods. In this work, we present Unified World Models (UWM), a framework that allows for leveraging both video and action data for policy learning. Specifically, a UWM integrates an action diffusion process and a video diffusion process within a unified transformer architecture, where independent diffusion timesteps govern each modality. By simply controlling each diffusion timestep, UWM can flexibly represent a policy, a forward dynamics, an inverse dynamics, and a video generator. Through simulated and real-world experiments, we show that: (1) UWM enables effective pretraining on large-scale multitask robot datasets with both dynamics and action predictions, resulting in more generalizable and robust policies than imitation learning, (2) UWM naturally facilitates learning from action-free video data through independent control of modality-specific diffusion timesteps, further improving the performance of finetuned policies. Our results suggest that UWM offers a promising step toward harnessing large, heterogeneous datasets for scalable robot learning, and provides a simple unification between the often disparate paradigms of imitation learning and world modeling. Videos and code are available at https://weirdlabuw.github.io/uwm/
Figure 1: Unified World Models integrates action and video diffusion in a unified transformer architecture controlled by modality-specific diffusion timesteps. The model can flexibly be trained on large robotics datasets and then flexibly perform a variety of different inferences at test time. Doing so naturally enables improved robustness and generalization for imitation learning.
@@ I Introduction
Imitation learning provides a simple and scalable way to imbue autonomous robots with complex behaviors using human demonstrations [6, 26, 11, 49]. Imitation learning via supervised learning, often referred to as behavior cloning (BC), has shown remarkable success due to the advent of powerful multimodal generative models such as diffusion [22] or flow-based models  [28]. With these methods, acquiring new behaviors amounts to collecting demonstrations and fitting a generative model to the action distributions given observations using a maximum likelihood (or closely related) objective. These models have shown robust and reliable behavior within the training distribution, but can be brittle when tasked beyond this distribution. A natural solution to synthesizing robust and generalizable controllers with imitation learning is to scale up the number of high-quality, on-robot demonstrations collected through robotic teleoperation. While achievable with considerable resources, this data scaling process is expensive and time-consuming. A natural question arises - are prevalent methodologies for imitation learning making maximal use of the available large-scale datasets?
@@
While imitation learning methods learn a mapping from states to optimal actions, they do not explicitly capture temporal dynamics that are naturally present in demonstration trajectories or videos. An alternative paradigm to direct imitation learning that can leverage such dynamics information is that of world modeling; learning approximate models of how the world changes over time. Commonly instantiated as predicting the future observations given current observations (and actions), world models can naturally be trained from large scale robotic datasets  [35, 46], but also from alternative sources of data such as uncurated “play” data [13] or even action-free data such as videos. A variety of world modeling techniques, such as video diffusion models  [23] or latent state-space modeling  [32], have shown impressive results on realistic generation of future frames. However, it is not yet clear how the ability of these world models to capture temporal dynamics can be brought to bear on improving the robustness and generalization of robotic controllers synthesized via imitation learning.
@@
In this work, we propose a new diffusion-based learning framework that unifies imitation learning and world modeling, incorporating knowledge of temporal dynamics gleaned from large robotic datasets into imitation learning policies. Our key insight is to integrate an action diffusion process and an image diffusion process into a single diffusion transformer model conditioned on independent diffusion timesteps. Leveraging a connection between diffusion noise at different timesteps of the forward diffusion noising process and partial masking, this allows for flexible sampling from a number of distributions simply by manipulating the diffusion timesteps independently at inference time. For example, to draw a sample from the policy, one can “mask out” the image diffusion process by fixing the diffusion timestep of image denoising to $T$, thereby marginalizing it. Similarly, one can draw a sample from the forward dynamics model by fixing the action diffusion timestep to 00, inferring next observations given current observations and “clean” actions. The same is also true of inverse dynamics models and unconditional video prediction models into the future. This yields a simple, unified diffusion model that can serve as a policy, dynamics model, video predictor or inverse model depending on the use case (Fig 1).
@@
Concretely, a UWM consists of a coupled score model that predicts action scores and future image scores, conditioned on the current image and corresponding diffusion timesteps that are separate for next-observation and action. During training, the timesteps are sampled independently at random, exposing the model to different combinations of image and action noises. During inference, UWMs enable flexible sampling from various distributions by manipulating the diffusion timesteps independently. In particular, a UWM can generate samples from (1) forward dynamics $p(o^{\prime}|o,a)$, (2) inverse dynamics $p(a|o,o^{\prime})$ (3) marginal action distribution (policy) $p(a|o)$, (4) marginal image distribution (video generative model) $p(o^{\prime}|o)$. We show that this learning framework leads to improved policies compared to standard imitation learning since, (1) the unified architecture enables feature sharing between action and pixels, resulting in additional supervision from the same data, (2) the model captures all combinations of marginal and conditional distributions, acquiring an understanding of the causal relationship between actions and images, (3) the model can learn from broader data modalities such as action-free videos.
@@
We demonstrate the effectiveness of UWM through a set of experiments across both simulation and real-world robotic manipulation tasks. We demonstrate that UWM is capable of extracting knowledge from multitask robotic datasets, and further leveraging action-free video data to improve its generalization to out-of-distribution conditions. These models are able to flexibly perform a variety of test-time inference, while retaining strong performance of both policy and dynamics prediction. We show that while conceptually simple, a number of careful design decisions must be made to enable strong performance of the UWM model. Through this investigation of UWM, we take a step towards bridging the gap between policies and world models for robot learning.
@@ II Preliminaries
Unified world models build on the framework of denoising diffusion models  [22], and their application to problems in robotic control  [11].
II-A Denoising Diffusion Models
@@
$q(x_{t}\mid x_{t-1})=\mathcal{N}\left(x_{t}\mid\sqrt{1-\beta_{t}}\,x_{t-1},\,\beta_{t}\,\mathbf{I}\right),\quad t=1,\dots,T.$
After $T$ steps, $x_{T}$ is nearly an isotropic Gaussian [22]. The corresponding reverse process aims to map $x_{T}$ back to a clean sample $x_{0}$ from the data distribution by iteratively denoising. While the exact reverse conditional $q(x_{t-1}\mid x_{t})$ is generally intractable, one learns a parametrized approximation,
@@
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\epsilon_{x}$ is the Gaussian noise injected into $x_{t}$. To approximate this conditional expectation $\mathbb{E}\left[\epsilon_{x}|x_{t}\right]$, DDPMs are trained using a variant of denoising score matching, which approximates this using a neural network $s_{\theta}$
@@
Given this reverse diffusion model, samples can approximately be drawn from the data distribution using a simple “denoising procedure”. Starting with a sample $x_{T}\sim\mathcal{N}(0,\,\mathbf{I})$ drawn from Gaussian noise, new samples are iteratively drawn from $p_{\theta}(x_{t-1}\mid x_{t})$ until a “clean sample” $x_{0}$ is obtained. This procedure allows for the representation of complex multimodal distributions where performing MLE tractably is challenging. Extensions have been applied to robotic control as well  [11].
@@
While the above-mentioned generative modeling process is unconditional, diffusion models are naturally extended to conditional settings. Consider a setting where multivariate data $(x_{0},z_{0})\sim p(x_{0},z_{0})$ is available, and the conditional distribution $p(x_{0}|z_{0})$ must be modeled. In these settings, a bulk of the machinery from above can be reused, simply with an additional conditioning variable. The forward process remains identical, while the reverse process is modified as
@@ III-A Problem Setup
We build on typical sequential decision making settings, assuming access to a dataset of (observation, action, next-observation) pairs $\mathcal{D}_{\text{e}}=\{(o_{i},a_{i},o^{\prime}_{i})\}_{i=1}^{N}$ provided by an expert demonstrator. For the sake of exposition, we will assume that the environment is Markovian in observations $o$. In addition to this action-labeled dataset, we may also be provided an action-free dataset $\mathcal{D}_{\text{af}}=\{(o_{i},o_{i+1})\}_{i=1}^{M}$. The question becomes - how can we extract the most learning signal out of these datasets for synthesizing robot controllers?
In this context, several different models may be desired - (1) a policy $p(a|o)$ (often referred to as $\pi(a|o)$), that samples optimal actions to execute at a particular observation, (2) a dynamics model $p(o^{\prime}|o,a)$, that samples future observations, given a current observation and action, (3) an inverse model, $p(a|o,o^{\prime})$ that predicts what distribution of actions can transition between a current observation and a desired next observation, (4) a video-prediction model $p(o^{\prime}|o)$ that predicts marginal future observations given current ones. While these models have each seen use in different contexts, they are largely considered to be disparate fields of study. In this work, we show that these are many sides of the same dice; they can be unified into a single model to benefit each other.
@@ III-B Unified World Models via Coupled Video-Action Diffusion
Figure 2: Unified World Model training and policy inference pipeline. The left panel shows UWM pretraining on robot trajectories with actions and co-training on action-free videos by masking out actions using diffusion timesteps. The right panel illustrates marginal and conditional inference modes, corresponding to the policy and the inverse dynamics.
The core idea of a UWM is to develop a single diffusion model that can be trained on samples from the joint distribution of data $p(o,a,o^{\prime})$ and used to flexibly perform inference for the policy $p(a|o)$, the dynamics model $p(o^{\prime}|o,a)$, the inverse model $p(a|o,o^{\prime})$ and a video prediction model $p(o^{\prime}|o)$, with simple modifications to test-time inference.
A unified world model instantiates a joint diffusion model that integrates next observation $o^{\prime}$ and action prediction $a$ into a single diffusion model conditioned on current observation $o$. This can naturally be done by parameterizing a joint noise prediction network $s_{\theta}(o,a_{t},o^{\prime}_{t},\,t)$ that approximates a conditional expectation over both action and next observation noise $\mathbb{E}\left[\epsilon_{a},\epsilon_{o^{\prime}}|o,a_{t},o^{\prime}_{t},\right]$, with $\epsilon_{a},\epsilon_{o^{\prime}}$ referring to noise on actions and next observations, and $t$ referring to the coupled timestep of the joint diffusion process. [^1] It is important to note that throughout this section $o$ will refer to the current observation, $o^{\prime}_{t}$ will refer to timestep $t$ in the diffusion process for the next observation, and $a_{t}$ will refer to timestep $t$ in the diffusion process for action. However, training such a joint noise prediction network [20] $s_{\theta}(o,a_{t},o^{\prime}_{t},\,t)$ does not accomplish flexible inference since it can only sample from the joint distribution of $(a,o^{\prime})$.
@@
A UWM models a coupled noise prediction network $s_{\theta}(o,a_{t_{a}},o^{\prime}_{t_{o^{\prime}}},t_{a},t_{o^{\prime}})$ that approximates a conditional expectation over noise $\mathbb{E}\left[\epsilon_{a},\epsilon_{o^{\prime}}|o,a_{t_{a}},o^{\prime}_{t_{o^{\prime}}}\right]$, with $\epsilon_{a},\epsilon_{o^{\prime}}$ referring to noise on next observations and actions, and $t_{a}$ and $t_{o^{\prime}}$ referring to the decoupled steps of the diffusion process with respect to actions and next-observations respectively. The ability to set diffusion timesteps independently allows for marginalization and conditioning of different variables. Fixing the timestep for either $t_{a}$ or $t_{o^{\prime}}$ to $T$ marginalizes the corresponding variable $a$, $o^{\prime}$, while setting the timestep to 00 performs conditioning. By setting timestep $t_{o^{\prime}}=T$, the joint model is approximating the expectation $\mathbb{E}\left[\epsilon_{a},\epsilon_{o^{\prime}},|o,a_{t_{a}},o^{\prime}_{T}\right]$. Since $o^{\prime}_{T}$ is approximately an isotropic Gaussian, this reduces to $\mathbb{E}\left[\epsilon_{a}|o,a_{t_{a}}\right]$, which represents a policy $p(a|o)$, thereby performing marginalization. Similarly, setting the timestep $t_{o^{\prime}}=0$ reduces the approximated distribution to $\mathbb{E}\left[\epsilon_{a}|o,a_{t_{a}},o^{\prime}\right]$ which corresponds to an inverse model $p(a|o,o^{\prime})$, thereby performing conditioning. Simply setting combinations of $t_{a}$ and $t_{o^{\prime}}$ allows for flexible inference of policies, dynamics models, inverse models, and video prediction from the same model!
This suggests a recipe for unified world modeling using a simple modification to the standard denoising objective [22]. To train a joint noise prediction diffusion model $(\epsilon^{\theta}_{a},\epsilon^{\theta}_{o^{\prime}})=s_{\theta}(o^{\prime}_{t_{o^{\prime}}},a_{t_{a}},o,t_{a},t_{o^{\prime}})$, we independently sample action timestep $t_{a}$ and next observation timestep $t_{o^{\prime}}$, draw noisy action and next-observation samples from their respective distributions, and train the coupled conditional score model conditioned on the current observation with a standard denoising objective across both next-observations and actions:
@@
where $w_{a}$ and $w_{o^{\prime}}$ are weights chosen to trade off between the action prediction and next-observation prediction objectives. Intuitively, this training paradigm exposes the model to all combinations of noise levels of the modalities, which enables flexible sampling at inference. We can use the trained model to flexibly draw samples from various distributions by simply controlling the time-steps $t_{a}$ and $t_{o^{\prime}}$ during inference, as follows:
@@
    Policy To sample from the policy $p(a|o)$, we must marginalize out the next observation $o^{\prime}$. To do so, we can simply set $t_{o^{\prime}}=T$ and perform denoising steps to synthesize an action $a$ conditioned on current observation $o$. We perform the reverse diffusion process on actions going from $t_{a}=T,\dots,1$, with frozen next observation timestep (at full masking) $t_{o^{\prime}}=T$, and $a_{T}\sim\mathcal{N}(0,I)$, $o^{\prime}_{T}\sim\mathcal{N}(0,I)$:
@@
    While this is simply performing action diffusion, it benefits from the temporal dynamics grounding obtained from the other modes.
@@
    Video Prediction Model To sample from the video prediction model $p(o^{\prime}|o)$, we must marginalize out the action $a$. We can simply set $t_{a}=T$ and perform denoising steps to synthesize the next observation $o^{\prime}$ conditioned on current observation $o$. We perform the reverse diffusion process on next observations going from $t_{o^{\prime}}=T,\dots,1$, with frozen action timestep (at full masking) $t_{o^{\prime}}=T$, and $a_{T}\sim\mathcal{N}(0,I)$, $o^{\prime}_{T}\sim\mathcal{N}(0,I)$:
@@
    Forward Dynamics In order to use the UWM as a forward dynamics model $p(o^{\prime}|o,a)$, we must condition on a particular executed action $a$. To do so, we can simply set $t_{a}=0$ and denoise in order to synthesize the next observation $o^{\prime}$ conditioned on current observation $o$ and the particular action $a$. We perform the reverse diffusion process on next observations going from $t_{o^{\prime}}=T,\dots,1$, with frozen action timestep (at no masking) $t_{a}=0$, and $o^{\prime}_{T}\sim\mathcal{N}(0,I)$:
@@
    Inverse Dynamics Lastly, to sample from the inverse dynamics model $p(a|o,o^{\prime})$, we must condition on a particular next observation $o^{\prime}$. To do so, we can simply set $t_{o^{\prime}}=0$ and perform denoising steps to synthesize the action $a$ conditioned on current observation $o$ and the particular next observation $o^{\prime}$. Specifically, we perform the reverse diffusion process on next observations going from $t_{a}=T,\dots,1$, with frozen next observation timestep (at no masking) $t_{o^{\prime}}=0$, and $a_{T}\sim\mathcal{N}(0,I)$:
@@
This simple modification to the standard diffusion training paradigm allows a single model to be trained, benefiting from feature sharing between different models of action and future observation prediction. This model can then be flexibly used for inference with just a simple of choice of time-step, making it a versatile, general-purpose decision-making model.

### v3

@@ Abstract
Imitation learning has emerged as a promising approach towards building generalist robots. However, scaling imitation learning for large robot foundation models remains challenging due to its reliance on high-quality expert demonstrations. Meanwhile, large amounts of video data depicting a wide range of environments and diverse behaviors are readily available. This data provides a rich source of information about real-world dynamics and agent-environment interactions. Leveraging this data directly for imitation learning, however, has proven difficult due to the lack of action annotation. In this work, we present Unified World Models (UWM), a framework that allows for leveraging both video and action data for policy learning. Specifically, a UWM integrates an action diffusion process and a video diffusion process within a unified transformer architecture, where independent diffusion timesteps govern each modality. By controlling each diffusion timestep, UWM can flexibly represent a policy, a forward dynamics, an inverse dynamics, and a video generator. Through simulated and real-world experiments, we show that: (1) UWM enables effective pretraining on large-scale multitask robot datasets with both dynamics and action predictions, resulting in more generalizable and robust policies than imitation learning, (2) UWM naturally facilitates learning from action-free video data through independent control of modality-specific diffusion timesteps, further improving the performance of finetuned policies. Our results suggest that UWM offers a promising step toward harnessing large, heterogeneous datasets for scalable robot learning, and provides a simple unification between the often disparate paradigms of imitation learning and world modeling. Videos and code are available at https://weirdlabuw.github.io/uwm/
Figure 1: Unified World Models integrates action and video diffusion in a unified transformer architecture controlled by modality-specific diffusion timesteps. The model can be trained on large robotics datasets and then flexibly perform a variety of different inferences at test time. This enables improved robustness and generalization for imitation learning.
@@ I Introduction
Imitation learning provides a simple and scalable way to imbue autonomous robots with complex behaviors using human demonstrations [6, 26, 11, 49]. Imitation learning via supervised learning, often referred to as behavior cloning (BC), has shown remarkable success due to the advent of powerful multimodal generative models such as diffusion [22] or flow-based models  [28]. With these methods, acquiring new behaviors amounts to collecting demonstrations and fitting a generative model to the action distributions given observations. However, despite showing robust and reliable behavior within the training distribution, these methods can be brittle when tasked beyond this distribution. A natural solution to synthesizing robust and generalizable controllers with imitation learning is to scale up the number of high-quality, on-robot demonstrations collected through robotic teleoperation. But this data scaling process is expensive and time-consuming. A natural question arises - are prevalent methodologies for imitation learning making maximal use of the available large-scale datasets?
@@
While imitation learning methods learn a mapping from states to optimal actions, they do not explicitly capture temporal dynamics that are naturally present in demonstration trajectories or videos. An alternative paradigm that can leverage such dynamics information is that of world modeling: learning approximate models of how the world changes over time. Commonly instantiated as predicting the future observations given current observations (and actions), world models can be trained from large scale robotic datasets  [35, 46], but also from alternative sources of data such as uncurated “play” data [13] or even action-free data such as videos. A variety of world modeling techniques, such as video diffusion models  [23] or latent state-space modeling  [32], have shown impressive results on realistic generation of future frames. However, it is not yet clear how the ability of these world models to capture temporal dynamics can be brought to bear on improving the robustness and generalization of robotic controllers synthesized via imitation learning.
@@
In this work, we propose a new diffusion-based learning framework that unifies imitation learning and world modeling, incorporating knowledge of temporal dynamics gleaned from large robotic datasets into imitation learning policies. Our key insight is to integrate an action diffusion process and an image diffusion process into a single diffusion transformer model conditioned on independent diffusion timesteps. Leveraging a connection between diffusion noise at different timesteps of the forward diffusion noising process and partial masking, this allows for flexible sampling from a number of distributions simply by manipulating the diffusion timesteps independently at inference time. For example, to draw a sample from the policy, one can “mask out” the image diffusion process by fixing the image diffusion timestep to $T$. Similarly, one can sample from the forward dynamics model by fixing the action diffusion timestep to 00, inferring next observations given current observations and “clean” actions. Same holds for inverse dynamics models and unconditional video prediction models into the future. This yields a simple, unified diffusion model that can serve as a policy, dynamics model, video predictor or inverse model (Fig 1).
@@
Concretely, a UWM consists of a coupled score model that predicts action scores and future image scores, conditioned on the current image and separate diffusion timesteps for action and future image. During training, the timesteps are sampled independently at random, exposing the model to different combinations of action and image noises. During inference, UWM enables flexible sampling from various distributions by manipulating the diffusion timesteps independently. In particular, a UWM can generate samples from (1) forward dynamics, (2) inverse dynamics (3) marginal action distribution (policy), (4) marginal image distribution (video generative model). We show that this learning framework leads to improved policies compared to standard imitation learning since, (1) the unified architecture enables feature sharing between actions and pixels, resulting in additional supervision from the same data, (2) the model captures all combinations of marginal and conditional distributions, acquiring an understanding of the causal relationship between actions and images, (3) the model can learn from broader data modalities such as action-free videos.
@@
We demonstrate the effectiveness of UWM through a set of experiments across both simulation and real-world robotic manipulation tasks. We show that UWM is capable of extracting knowledge from multitask robotic datasets, and further leveraging action-free video data to improve its generalization to out-of-distribution conditions. These models are able to flexibly perform a variety of test-time inference, while retaining strong performance of both policy and dynamics prediction. Through this investigation of UWM, we take a step towards bridging the gap between policies and world models for robot learning.
@@ II Preliminaries
Unified world models build on the framework of denoising diffusion probabilistic models  [22] and their application to problems in robotic control  [11].
II-A Diffusion Models
@@
$q(x_{t}\mid x_{t-1})=\mathcal{N}\left(x_{t}\mid\sqrt{1-\beta_{t}}\,x_{t-1},\,\beta_{t}\,I\right),\quad t=1,\dots,T.$
After $T$ steps, $x_{T}$ is nearly an isotropic Gaussian [22]. The corresponding reverse process aims to map $x_{T}$ back to a clean sample $x_{0}$ from the data distribution by iteratively denoising. While the exact reverse conditional $q(x_{t-1}\mid x_{t})$ is generally intractable, one can learn a parametrized approximation,
@@
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\epsilon_{x}$ is the Gaussian noise injected into $x_{t}$. To approximate this conditional expectation $\mathbb{E}\left[\epsilon_{x}|x_{t}\right]$, DDPMs train a neural network $s_{\theta}$ using a variant of denoising score matching:
@@
Given this reverse diffusion model, samples can approximately be drawn from the data distribution using a simple “denoising procedure”. Starting with a sample $x_{T}\sim\mathcal{N}(0,\,I)$ drawn from Gaussian noise, new samples are iteratively drawn from $p_{\theta}(x_{t-1}\mid x_{t})$ until a “clean sample” $x_{0}$ is obtained. This procedure allows for the representation of complex multimodal distributions where performing MLE tractably is challenging.
@@
While the above-mentioned generative modeling process is unconditional, diffusion models can be naturally extended to conditional settings. Consider a setting where multivariate data $(x_{0},z_{0})\sim p(x_{0},z_{0})$ is available, and the conditional distribution $p(x_{0}|z_{0})$ must be modeled. In these settings, a bulk of the machinery from above can be reused, simply with an additional conditioning variable. The forward process remains identical, while the reverse process is modified as
@@ III-A Problem Setup
We build on typical sequential decision making settings, assuming access to a dataset of (observation, action, next-observation) pairs $\mathcal{D}_{\text{e}}=\{(o_{i},a_{i},o^{\prime}_{i})\}_{i=1}^{N}$ provided by an expert demonstrator. For the sake of exposition, we will assume that the environment is Markovian in observations $o$. In addition to this action-labeled dataset, we may also have access to an action-free dataset $\mathcal{D}_{\text{af}}=\{(o_{i},o_{i+1})\}_{i=1}^{M}$. The question becomes - how can we extract the most learning signal out of these datasets for synthesizing robot controllers?
In this context, several different models may be desired: (1) a policy $p(a|o)$ (often referred to as $\pi(a|o)$) that samples optimal actions to execute at a particular observation, (2) a dynamics model $p(o^{\prime}|o,a)$ that samples future observations, given a current observation and action, (3) an inverse model $p(a|o,o^{\prime})$ that predicts what distribution of actions can transition between a current observation and a desired next observation, and (4) a video prediction model $p(o^{\prime}|o)$ that predicts marginal future observations given current ones. While these models have each seen use in different contexts, they are largely considered to be disparate fields of study. In this work, we show that these are many sides of the same dice; they can be unified into a single model to benefit each other.
@@ III-B Unified World Models via Coupled Video-Action Diffusion
Figure 2: Unified World Model training and inference pipeline. The left panel shows UWM pretraining on robot trajectories with actions and co-training on action-free videos by masking out actions using diffusion timesteps. The right panel illustrates marginal and conditional inference modes, corresponding to the policy and the inverse dynamics.
The core idea of a UWM is to develop a single diffusion model that can be trained on samples from the joint distribution of data $p(o,a,o^{\prime})$ and used to flexibly perform inference for the policy $p(a|o)$, the dynamics model $p(o^{\prime}|o,a)$, the inverse model $p(a|o,o^{\prime})$ and the video prediction model $p(o^{\prime}|o)$, with minimal modifications to test-time inference.
We start by instantiating a joint diffusion model that integrates next observation prediction $o^{\prime}$ and action prediction $a$ into a single diffusion model conditioned on current observation $o$. This can naturally be done by parameterizing a joint noise prediction network $s_{\theta}(o,a_{t},o^{\prime}_{t},\,t)$ that approximates a conditional expectation over both action and next observation noise $\mathbb{E}\left[\epsilon_{a},\epsilon_{o^{\prime}}|o,a_{t},o^{\prime}_{t},\right]$, with $\epsilon_{a},\epsilon_{o^{\prime}}$ referring to noise on actions and next observations, and $t$ referring to the coupled timestep of the joint diffusion process. [^1] It is important to note that throughout this section $o$ will refer to the current observation, $o^{\prime}_{t}$ will refer to timestep $t$ in the diffusion process for the next observation, and $a_{t}$ will refer to timestep $t$ in the diffusion process for action. However, training such a joint noise prediction network [20] $s_{\theta}(o,a_{t},o^{\prime}_{t},\,t)$ does not accomplish flexible inference since it can only sample from the joint distribution of $(a,o^{\prime})$.
@@
Concretely, a UWM models a coupled noise prediction network $s_{\theta}(o,a_{t_{a}},o^{\prime}_{t_{o^{\prime}}},t_{a},t_{o^{\prime}})$ that approximates a conditional expectation over noise $\mathbb{E}\left[\epsilon_{a},\epsilon_{o^{\prime}}|o,a_{t_{a}},o^{\prime}_{t_{o^{\prime}}}\right]$, with $\epsilon_{a},\epsilon_{o^{\prime}}$ referring to noise on next observations and actions, and $t_{a}$ and $t_{o^{\prime}}$ referring to the decoupled steps of the diffusion process with respect to actions and next-observations respectively. The ability to set diffusion timesteps independently allows for marginalization and conditioning of different variables. Fixing the timestep for either $t_{a}$ or $t_{o^{\prime}}$ to $T$ marginalizes the corresponding variable $a$, $o^{\prime}$, while setting the timestep to 00 performs conditioning. By setting timestep $t_{o^{\prime}}=T$, the joint model is approximating the expectation $\mathbb{E}\left[\epsilon_{a},\epsilon_{o^{\prime}},|o,a_{t_{a}},o^{\prime}_{T}\right]$. Since $o^{\prime}_{T}$ is approximately an isotropic Gaussian, this reduces to $\mathbb{E}\left[\epsilon_{a}|o,a_{t_{a}}\right]$, which represents a policy $p(a|o)$, thereby performing marginalization. Similarly, setting the timestep $t_{o^{\prime}}=0$ reduces the approximated distribution to $\mathbb{E}\left[\epsilon_{a}|o,a_{t_{a}},o^{\prime}\right]$ which corresponds to an inverse model $p(a|o,o^{\prime})$, thereby performing conditioning. Simply setting combinations of $t_{a}$ and $t_{o^{\prime}}$ allows for flexible inference of policies, dynamics models, inverse models, and video prediction from the same model.
This suggests a training recipe using a simple modification to the standard denoising objective [22]. To train a joint noise prediction diffusion model $(\epsilon^{\theta}_{a},\epsilon^{\theta}_{o^{\prime}})=s_{\theta}(o^{\prime}_{t_{o^{\prime}}},a_{t_{a}},o,t_{a},t_{o^{\prime}})$, we independently sample action timestep $t_{a}$ and next observation timestep $t_{o^{\prime}}$, draw noisy action and next-observation samples from their respective distributions, and train the coupled conditional score model conditioned on the current observation with a standard denoising objective across both actions and next-observations:
@@
where $w_{a}$ and $w_{o^{\prime}}$ are weights chosen to trade off between the action prediction and next-observation prediction objectives. Intuitively, this training paradigm exposes the model to all combinations of noise levels of the modalities. At inference, we can flexibly draw samples from various distributions by controlling the timesteps $t_{a}$ and $t_{o^{\prime}}$ as follows:
@@
    Policy To sample from the policy $p(a|o)$, we marginalize out the next observation $o^{\prime}$ by setting $t_{o^{\prime}}=T$ and $o^{\prime}_{T}\sim\mathcal{N}(0,I)$. We perform the reverse diffusion process on actions going from $t_{a}=T,\dots,1$ with $a_{T}\sim\mathcal{N}(0,I)$:
@@
    Video Prediction Model To sample from the video prediction model $p(o^{\prime}|o)$, we marginalize out the action $a$ by setting $t_{a}=T$ and $a_{T}\sim\mathcal{N}(0,I)$. We perform the reverse diffusion process on next observations going from $t_{o^{\prime}}=T,\dots,1$ with $o^{\prime}_{T}\sim\mathcal{N}(0,I)$:
@@
    Forward Dynamics To sample from the forward dynamics model $p(o^{\prime}|o,a)$, we condition on a particular action $a$ by setting $t_{a}=0$ and $a_{0}=a$. We perform the reverse diffusion process on next observations going from $t_{o^{\prime}}=T,\dots,1$ with $o^{\prime}_{T}\sim\mathcal{N}(0,I)$:
@@
    Inverse Dynamics To sample from the inverse dynamics model $p(a|o,o^{\prime})$, we condition on a particular next observation $o^{\prime}$ by setting $t_{o^{\prime}}=0$ and $o_{0}=o$. We perform the reverse diffusion process on actions going from $t_{a}=T,\dots,1$ with $a_{T}\sim\mathcal{N}(0,I)$:
@@
This simple modification to the standard diffusion training paradigm allows a single model to be trained, benefiting from feature sharing between different models of action and future observation prediction. This model can then be flexibly used for inference with just the choice of timesteps, making it a versatile, general-purpose decision-making model.